OpenStudy (anonymous):

Write an equation of a line with slope m = -7/8 and y-intercept (0,-5)

OpenStudy (anonymous):

(y+5)=(-7/8)(x-0) then 8y+7x+40 = 0

OpenStudy (anonymous):

Since you know the slope, now all you have to do is use this formula y-y1=m(x-x1) 0=x1 -5=y1

OpenStudy (anonymous):

These are my choices..

OpenStudy (anonymous):

8y+7x+40 = 0 -> y=-(7x+40)/8=(-7x/8)-5 ->x=-(8y+40)/7=(-8y/7)-(40/7)